A spring with a spring constant of 950 N/m is compressed 0.20 m. What speed can it give to a 1.5 kg ball when it is released?


Please answer using grade 11 Conservation of Energy and Momentum please!




1
Expert's answer
2020-05-11T20:06:13-0400

Given Data : spring constant K = 950 N/m x = 0.20 m, Ball Mass m = 1.5 Kg


Energy stored in spring = "\frac{1}{2} \times k \times x^{2}"


Energy stored in spring =  "\frac{1}{2} \times 950 \times 0.20^{2}"

 

= 19 N.m (J)


Potential energy stored in spring = Kinetic Energy released to Ball


19 =  "\frac{1}{2} \times m \times v^{2}"



19 =  "\frac{1}{2} \times 1.5 \times v^{2}"


V =  "\sqrt{25.33}"


V = 5.03 "\frac{m}{sec^{2}}"


Speed of Ball is 5.03 "\frac{m}{sec^{2}}"